OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (anonymous):

i think its y + 1 = 4(x + 3)

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

What is the domain of f(x)? -2 < x ≤ 3 -2 ≤ x < 3 -1 < y ≤ 3 -1 ≤ y < 3

jigglypuff314 (jigglypuff314):

domain has to do with x which is horizontally closed hole means "or equal to" open hole does not have the "or equal to" line

OpenStudy (anonymous):

is it -1 < y ≤ 3

OpenStudy (acxbox22):

try again

jigglypuff314 (jigglypuff314):

domain has to do with x

OpenStudy (acxbox22):

that is the range

OpenStudy (anonymous):

so then it would be -2 ≤ x < 3

jigglypuff314 (jigglypuff314):

closed hole at x = 3 so x ≤ 3

OpenStudy (anonymous):

ok

OpenStudy (acxbox22):

the other one

OpenStudy (anonymous):

so then its the oppisite way around which is-2 < x ≤ 3

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

If x represents time, the average rate of change of the function f(x) in the first three seconds is ___

jigglypuff314 (jigglypuff314):

average rate of change = slope so figure out the slope

OpenStudy (anonymous):

slope = 4/3??

OpenStudy (anonymous):

or 3/2 or 1 im not sure

jigglypuff314 (jigglypuff314):

mmm not quite? it's going down so negative, first of all then try making it out of points (1, 100) , (3, 0)

OpenStudy (anonymous):

5/4

jigglypuff314 (jigglypuff314):

errr do you have options?

OpenStudy (anonymous):

1

jigglypuff314 (jigglypuff314):

huh? well, I'm thinking the misconception is that the grid isn't by ones like across it goes 0, 0.5, 1, 1.5, 2, 2.5, ect while up it goes 0, 20, 40, 60, etc so if you look at the slope based on the numbers and not grids take the points (1, 100) and (3, 0) and find the slope :/

OpenStudy (anonymous):

100/7

jigglypuff314 (jigglypuff314):

mmm I'm getting -100/2

OpenStudy (anonymous):

i can kinda see were you got that

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } = \frac{ 0-100 }{ 3-1 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

here just help me on this one

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

jigglypuff314 (jigglypuff314):

slope-intercept form: y = mx + b when m = slope and b = y-intercept

OpenStudy (anonymous):

i got y=-3x+13

OpenStudy (anonymous):

-3x+13

jigglypuff314 (jigglypuff314):

try again? y = -3x + b -5 = -3(2) + b

OpenStudy (anonymous):

+11

OpenStudy (anonymous):

no it would be a negative right

jigglypuff314 (jigglypuff314):

-5 = -3(2) + b -5 = -6 + b +6 +6 1 = b

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (1, -5). y = -5 x = 1 y = 1 x = -5

jigglypuff314 (jigglypuff314):

horizontal line is y = (something) (x, y) = (1, -5)

OpenStudy (anonymous):

i think its x=1

OpenStudy (anonymous):

i thought horizontal line is x

jigglypuff314 (jigglypuff314):



OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so then it would be y=-5

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

y = -2x - 6 y = negative two thirdsx + 6 y = -negative two thirdsx + 6 y = 2x + 6

OpenStudy (anonymous):

i think its y= - 2/3+6

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

never mind i got it thanks for all the help

jigglypuff314 (jigglypuff314):

you're welcome :)